# Workload-Aware Incremental Reclustering in Cloud Data Warehouses

### v1

@@ 1. INTRODUCTION
For large tables with continuous data ingestion, manually maintaining the clustered state places a huge burden on database users to monitor and analyze DML patterns and workload distributions. Existing approaches, such as Qd-tree (Yang et al., 2020) and MDDL in Redshift (Ding et al., 2024), assume a stationary workload and learn an optimal layout for the entire table from historical queries. They are, however, suboptimal under continuous data ingestion and evolving workload patterns. Iceberg (7), Delta Lake (21; 20), and Databricks (18; 17) allow dynamically switching clustering keys to accommodate workload shifts, but the updated physical layout is only applied to newly ingested data. Snowflake (42) and Dremio (28) propose the data-driven incremental reclustering, where they develop static metrics to evaluate the clustering quality of each micro-partition and only recluster the ill-clustered ones. Although these approaches reduce the overall reclustering cost, they are unaware of workloads, resulting in unnecessary reclustering of rarely accessed data.
@@
We make four primary contributions in this paper. First, we propose a policy-key decoupled taxonomy and framework for reclustering. Second, we introduce the concept of boundary micro-partitions that dominate scan-set pruning efficiency, and prove a theoretical upper bound for the cost of reclustering all these boundary micro-partitions margin: under possible workload shifts and constraints. Third, we propose WAIR based on boundary micro-partitions that can automatically determine the reclustering timing and the micro-partition set to maximize overall cost reduction. Finally, we implement a workload-aware automatic reclustering service based on WAIR and demonstrate its advantages in cost, query performance, and robustness against diverse workloads over existing solutions.
@@ 2. PRELIMINARIES
Figure 1. Micro-partitions in Cloud Data Warehouses. A table is structured as multiple micro-partitions stored in the object storage layer. Each micro-partition comprises both metadata (e.g., zonemaps) and data blocks. During query execution, the metadata is used to prune irrelevant partitions.
@@
A defining feature of cloud data warehouses is the disaggregation of compute and storage for independent scaling (Zaharia et al., 2021; Barnhart et al., 2024; Dageville et al., 2016). For queries on large tables, avoiding full table scans is critical for achieving optimal analytical performance. A widely adopted approach is partitioning (Larson et al., 2013; Yang et al., 2020; Ding et al., 2021b). In cloud data warehouses such as Snowflake, tables are divided into micro-partitions [^1] Throughout this paper, the term “partition” typically refers to micro-partitions, unless otherwise specified. (Dageville et al., 2016), a concept similar to “row groups” in columnar data formats like Parquet (10) and ORC (9). As illustrated in Figure 1, incoming data is automatically partitioned into immutable files (i.e., micro-partitions) according to its natural ingestion order. These micro-partitions reside in cloud object storage (e.g., AWS S3, Azure Blob Storage) and serve as fundamental units for both data retrieval and pruning. Each micro-partition is maintained at a fixed size, typically a few tens of megabytes (e.g., 32MB) (Durner et al., 2023).
@@ 2.1. Benefits and Challenges of Clustering
Figure 2. Benefits of Clustering. The query predicated on Age must scan all partitions in a naturally ordered table, while clustering by Age allows pruning all but the first partition.
@@ 2.2. Related Work
The term “partitioning” is overloaded in database literature. A significant body of work focuses on horizontal partitioning (or sharding) for distributed databases (e.g., Azure Synapse SQL (J. Aguilar-Saborit and R. Ramakrishnan (2020); 26)) or load balancing in streaming systems (e.g., Dalton (Zapridou et al., 2022)). The primary goal of these techniques is to minimize cross-node traffic and distribute compute load. In contrast, modern cloud data warehouses adopt micro-partitioning within the cloud storage layer for data skipping. This relies on lightweight metadata (e.g., zone maps) (Ziauddin et al., 2017; Zimmerer et al., 2025) and space-filling curves (e.g., Z-ordering (Morton, 1966)) to prune irrelevant micro-partitions during execution. In this paper, we target this micro-partition level optimization to facilitate effective pruning, which is orthogonal to and can coexist with prior works.
Policy Key Fixed Dynamic- Manual Dynamic- Workload
Full Table legacy Qd-tree (Yang et al., 2020) Redshift (4)
New Data Iceberg (8) Delta Lake (22) Databricks (19)
margin: Incremental- Data-Driven Snowflake (43) Dremio (29)
Incremental- Workload-Aware WAIR (ours)
@@
Iceberg supports in-place table evolution, where newly ingested data follow the updated partition layout or sort order (7). Delta Lake organizes files into stable “ZCubesZCubes,” each produced by a single OPTIMIZE job. During every optimization cycle, it clusters newly ingested files together with any undersized, partial $ZCubes$ left behind by DELETE operations (21; 20). Databricks extends Delta Lake by analyzing query workloads to identify the most effective clustering columns, then applies those keys to subsequent ingests (18; 17). These approaches have two main drawbacks: 1) Existing historical data remains untouched, and 2) independent “sorted runs” across batches degrade global clustering quality over time.
@@
Snowflake evaluates a table’s clustering quality by the overlapping metrics (42). Using partition metadata, it counts the number of micro-partitions in the table whose key ranges overlap with a given micro-partition and a given point (see Figure 3). These metrics averaged across the table reflect its overall clustering quality. Dremio uses the same metric, where only a small subset of files with the highest overlapping depth within their key ranges are reclustered at optimization cycles, until the table satisfies a user-defined overlapping depth threshold (28). While more efficient than full repartitioning, these data-driven techniques: 1) involves potential expensive scans to compute metrics, 2) rely on manual parameter tuning; 3) ignore query patterns, resulting in wasteful reclustering of rarely accessed historical data.
@@ 3. Boundary Micro-Partitions
A key observation behind incremental reclustering is that most of the query performance benefits comes from reclustering only a strategic subset of partitions. In this section, we introduce the concept of boundary partitions and explain why they are critical for pruning efficiency. We then propose a simple greedy reclustering algorithm based on boundary partitions with a theoretical analysis of its amortized cost.
@@ 3.1. The Reclustering Workflow
Figure 4 shows the typical workflow in a cloud data warehouse. An initial pool of micro-partitions exists in cloud object storage and is clustered based on certain clustering keys. A clustering key is the column(s) used to sort a micro-partition. Without loss of generality, we assume a clustering key contains a single column. New data is periodically ingested into database tables as micro-partitions. These incoming micro-partitions are clustered in their natural ingestion order, which is typically different from the clustering-key order.
@@ 3.2. Micro-Partitions on the Query Boundaries
A basic assumption of workload-aware algorithms is that future queries follow a similar statistical pattern as the historical ones. As depicted in Figure 5, we define boundary micro-partitions as those that partially overlap with at least one already-queried range. Each range query generates two sets of boundary partitions that overlap either the left or right border of its range predicate. Micro-partitions that fall completely inside or outside any query ranges do not require reclustering because all the records in them are either included or excluded from query results. Boundary micro-partitions, on the other hand, are critical to pruning efficiency because scanning these micro-partitions leads to retrieving unnecessary rows and therefore wastes I/O bandwidth. Algorithm 1 presents a basic greedy algorithm that reclusters all boundary micro-partitions after each query. We then provide a theoretical analysis to show that this algorithm incurs at most logarithmic additional cost.
@@
For all but $O(\log q)$ matched input-output pairs,
$\phi(c_{j},d_{j})\leq\phi(a_{i},b_{i})-4$.
@@ 4.2. Cost-based Trade-off
$P_{i}]=sort(\{\Delta\hat{cost}_{q}(P),P\in\mathcal{P}\})$
for $cut\in\{1,2,\dots,m\}$ do
@@
$cut\leftarrow\mathop{\arg\min}_{cut}\Delta\hat{cost} \triangleright$ the smaller the better
@@
   current $\Delta cost+\hat{cost}_{r}([P_{i}]_{cut})<cost\_limit$ then
  Recluster($P_{i}]_{cut}$)
Additionally, we dynamically adjust the sliding window size based on its prediction accuracy to adapt to the evolution of the query workload and the data distribution. After completing a window $W$ of queries, we compare the actual cost reduction $\Delta cost_{q}=\sum_{Q\in W}\Delta cost_{q}(Q)$ against the predicted cost reduction $\Delta\hat{cost}_{q}=\sum_{Q\in W}\Delta\hat{cost}_{q}(Q)$. If $\Delta cost_{q}>\Delta\hat{cost}_{q}$, it indicates that the workload is relatively stable, and we therefore double the window size to capture a longer query history. Otherwise, we halve the window size to focus more narrowly on the most recent workload shifts.
@@ 4.3. Cost Model Extensions
where $|\Delta\hat{pruned}(\mathcal{P}\mid Q)|$ denotes the estimated increase in skipped micro-partitions for $Q$ after reclustering $\mathcal{P}$. By tuning $\beta$, users can express specific exchange rates between budget and performance.
@@ 4.4. Hybrid Layout
To understand how each micro-partition might benefit from reclustering, we decompose its aggregate estimated savings into per-column components. We assume that each predicated column contributes equally in a query. Specifically, for each selected micro-partition $P$, we distribute its aggregated savings across all columns that appear in a query’s predicate within the sliding window: $s_{P}=\bigl(s_{P}^{(1)},s_{P}^{(2)},\dots,s_{P}^{(d)}\bigr),$ where $d$ is the number of attributes, $s_{P}^{(i)}$ is the share of $P$’s projected benefit attributed to column $i$. Each savings vector is then plotted in a high-dimensional space to create the micro-partition’s “savings signature”, as shown in Figure 8.
@@
Hybrid Layout for Clusters
  • Single-column dominance. If a micro-partition’s projected savings are contributed almost entirely by a single column, reclustering using the column as the clustering key is sufficient. Micro-partitions within the same cluster are grouped and sorted together by its dominant column.
@@
The candidate micro-partitions are therefore divided into mutually exclusive groups, each assigned the physical layout that promises the greatest benefit. This strategy produces a hybrid table organization: different regions of the data are clustered with different keys, yet every reclustering decision is still guided by the same workload-aware cost model described earlier. Micro-partitions can thus shift between groups as the workload evolves, allowing the table layout to adapt over time.
@@ 5. RECLUSTERING AS A SERVICE
Having presented the WAIR algorithm, we now introduce the Automatic Reclustering Service (ARS), a scalable reclustering framework which autonomously manages background operations to preserve table clustering quality. ARS leverages execution statistics with minimal compute overhead to decide the scope and timing of reclustering. ARS moves all reclustering work off the query-execution critical path, ensuring zero disruption to production workloads. In this section, we describe how to embed the WAIR algorithm into ARS with minimal engineering effort and margin: full compatibility on top of distributed cloud data warehouses. We present the practical feasibility of ARS and implementation details to be evaluated and margin: outline an overview of pruning opportunities in general.
@@ 5.1. Service Framework
Integrating the reclustering framework into an existing cloud data warehouse additionally requires: 1) an (often already in place) metadata service, and 2) a dedicated reclustering job executor.
@@
Metadata is tracked and preserved during data ingestion and query execution asynchronously for further utilization. margin: In distributed systems, the metadata service is typically a standalone component. It should integrate seamlessly with a data warehouse’s existing metadata infrastructure as a drop-in module.
Reclustering Executor. After specifying the partition set to be reclustered and corresponding layouts, the reclustering executor retrieves these partitions from storage layer, loads their data into local memory, and reorganize them into new partitions. These clustered partitions are then persisted back to the object storage.
@@
Reclustering operations are designed to be non-blocking in distributed systems. When a reclustering operation completes, ARS publishes a new table snapshot that references the current set of partitions. Because reclustering reorganizes data without changing any row values, queries that began before completion continue to read the prior snapshot along with any newly ingested partitions. New queries started after completion read the reclustered partitions referenced by the latest, updated snapshot. Once no queries reference an older snapshot, ARS safely removes its obsolete partitions during garbage collection. The table snapshot chain incurs minimal storage overhead and allows new queries to benefit immediately once reclustering completes. Reclustering tasks run on dynamically provisioned compute nodes, entirely outside the query execution critical path. margin: In cloud data warehouses with a cache layer, reclustering can naturally benefit from scan sharing as it requires scanning complete micro-partition data.

### v2

@@ Abstract.
cloud data warehouses; analytical query processing; micro-partitions; zonemap pruning; incremental table reclustering; workload-aware optimization; cost-based optimization
copyright: ccjournal: PACMMODjournalyear: 2026journalvolume: 4journalnumber: 3 (SIGMOD)article: 250publicationmonth: 6doi: 10.1145/3802127ccs: Information systems Data warehousesccs: Information systems Cloud based storageccs: Information systems Database query processingccs: Information systems Autonomous database administration
@@ 1. INTRODUCTION
For large tables with continuous data ingestion, manually maintaining the clustered state places a huge burden on database users to monitor and analyze DML patterns and workload distributions. Existing approaches, such as Qd-tree (Yang et al., 2020) and MDDL in Redshift (Ding et al., 2024), assume a stationary workload and learn an optimal layout for the entire table from historical queries. They are, however, suboptimal under continuous data ingestion and evolving workload patterns. Iceberg (8), Delta Lake (22; 21), and Databricks (19; 18) allow dynamically switching clustering keys to accommodate workload shifts, but the updated physical layout is only applied to newly ingested data. Snowflake (43) and Dremio (29) propose the data-driven incremental reclustering, where they develop static metrics to evaluate the clustering quality of each micro-partition and only recluster the ill-clustered ones. Although these approaches reduce the overall reclustering cost, they are unaware of workloads, resulting in unnecessary reclustering of rarely accessed data.
@@
We make four primary contributions in this paper. First, we propose a policy-key decoupled taxonomy and framework for reclustering. Second, we introduce the concept of boundary micro-partitions that dominate scan-set pruning efficiency, and prove a theoretical upper bound for the cost of reclustering all these boundary micro-partitions. Third, we propose WAIR based on boundary micro-partitions that can automatically determine the reclustering timing and the micro-partition set to maximize overall cost reduction. Finally, we implement a workload-aware automatic reclustering service based on WAIR and demonstrate its advantages in cost, query performance, and robustness against diverse workloads over existing solutions.
@@ 2. PRELIMINARIES
Figure 1. Micro-partitions in Cloud Data Warehouses. A table is structured as multiple micro-partitions stored in the object storage layer. Each micro-partition comprises both metadata (e.g., zonemaps) and data blocks. During query execution, the metadata is used to prune irrelevant micro-partitions.
@@
A defining feature of cloud data warehouses is the disaggregation of compute and storage for independent scaling (Zaharia et al., 2021; Barnhart et al., 2024; Dageville et al., 2016). For queries on large tables, avoiding full table scans is critical for achieving optimal analytical performance. A widely adopted approach is partitioning (Larson et al., 2013; Yang et al., 2020; Ding et al., 2021b). In cloud data warehouses such as Snowflake, tables are divided into micro-partitions [^1] Throughout this paper, the term “partition” typically refers to micro-partitions, unless otherwise specified. (Dageville et al., 2016), a concept similar to “row groups” in columnar data formats like Parquet (11) and ORC (10). As illustrated in Figure 1, incoming data is automatically partitioned into immutable files (i.e., micro-partitions) according to its natural ingestion order. These micro-partitions reside in cloud object storage (e.g., AWS S3, Azure Blob Storage) and serve as fundamental units for both data retrieval and pruning during query processing. Each micro-partition is maintained at a fixed size, typically a few tens of megabytes in practice (e.g., 32MB) (Durner et al., 2023).
@@ 2.1. Benefits and Challenges of Clustering
Figure 2. Benefits of Clustering. The query predicated on Age must scan all micro-partitions in a naturally ordered table, while clustering by Age allows pruning all but the first micro-partition.
@@ 2.2. Related Work
Policy Key Fixed Dynamic (Manual) Dynamic (Workload)
Full Table legacies Qd-tree (Yang et al., 2020), Redshift (4)
New Data Iceberg (9), Delta Lake (23) Databricks (20)
margin: Incremental (Data-Driven) Snowflake (44), Dremio (30)
Incremental (Workload-Aware) WAIR (ours)
@@
The term “partitioning” is overloaded in database literature. A significant body of work focuses on horizontal partitioning (or sharding) for distributed databases (e.g., Azure Synapse SQL (J. Aguilar-Saborit and R. Ramakrishnan (2020); 27)) or load balancing in streaming systems (e.g., Dalton (Zapridou et al., 2022)). The primary goal of these techniques is to minimize cross-node traffic and distribute compute load. In contrast, modern cloud data warehouses adopt micro-partitioning within the cloud storage layer for data skipping. This relies on lightweight metadata (e.g., zone maps) (Ziauddin et al., 2017; Zimmerer et al., 2025) and space-filling curves (e.g., Z-ordering (Morton, 1966)) to prune irrelevant micro-partitions during execution. In this paper, we target this micro-partition level optimization to facilitate effective pruning, which is orthogonal to and can coexist with prior works.
@@
Iceberg supports in-place table evolution, where newly ingested data follow the updated partition layout or sort order (8). Delta Lake organizes files into stable “ZCubesZCubes,” each produced by a single OPTIMIZE job. During every optimization cycle, it clusters newly ingested files together with any undersized, partial $ZCubes$ left behind by DELETE operations (22; 21). Databricks extends Delta Lake by analyzing query workloads to identify the most effective clustering columns, then applies those keys to subsequent ingests (19; 18). These approaches have two main drawbacks: 1) Existing historical data remains untouched, and 2) independent “sorted runs” across batches degrade global clustering quality over time.
Snowflake evaluates a table’s clustering quality by the overlapping metrics (43). Using partition metadata, it counts the number of micro-partitions in the table whose key ranges overlap with a given micro-partition and a given point (see Figure 3). These metrics averaged across the table reflect its overall clustering quality. Dremio uses the same metric, where only a small subset of files with the highest overlapping depth within their key ranges are reclustered at optimization cycles, until the table satisfies a user-defined overlapping depth threshold (29). While more efficient than full repartitioning, these data-driven techniques: 1) involve potential expensive scans to compute metrics, 2) rely on manual parameter tuning; 3) ignore query patterns, resulting in wasteful reclustering of rarely accessed historical data.
@@ 3. Boundary Micro-Partitions
A key observation behind incremental reclustering is that most of the query performance benefits come from reclustering only a strategic subset of partitions. In this section, we introduce the concept of boundary partitions and explain why they are critical for pruning efficiency. We then propose a simple greedy reclustering algorithm based on boundary partitions with a theoretical analysis of its amortized cost.
@@ 3.1. The Reclustering Workflow
Figure 4 shows the typical workflow in a cloud data warehouse. An initial pool of micro-partitions exists in cloud object storage and is clustered based on certain clustering keys. A clustering key is the column(s) used to sort a micro-partition. For the theoretical analysis in this section, we assume a clustering key contains a single column. New data is periodically ingested into database tables as micro-partitions. These incoming micro-partitions are clustered in their natural ingestion order, which is typically different from the clustering-key order.
@@ 3.2. Micro-Partitions on the Query Boundaries
A basic assumption of workload-aware algorithms is that future queries follow a similar statistical pattern as the historical ones. As depicted in Figure 5, we define boundary micro-partitions as those that partially overlap with at least one already-queried range. Each range query generates two sets of boundary partitions that overlap either the left or right border of its range predicate. Micro-partitions that fall completely inside or outside any query ranges do not require reclustering because all the records in them are either included or excluded from query results. Boundary micro-partitions, on the other hand, are critical to pruning efficiency because scanning these micro-partitions leads to retrieving unnecessary rows and therefore wastes I/O bandwidth during query execution. Algorithm 1 presents a basic greedy algorithm that reclusters all boundary micro-partitions after each query. We then provide a theoretical analysis to show that this algorithm incurs at most logarithmic additional cost.
@@
For all but $O(\log q)$ matched input-output pairs, $\phi(c_{j},d_{j})\leq\phi(a_{i},b_{i})-4$.
@@ 4.2. Cost-based Trade-off
$P_{i}]=sort_{desc}(\{\Delta\hat{cost}_{q}(P),P\in\mathcal{P}\})$
for $cut\in\{1,2,\dots,\lvert\mathcal{P}\rvert\}$ do
@@
$cut^{*}\leftarrow\mathop{\arg\min}_{cut}\Delta\hat{cost} \triangleright$ the smaller the better
@@
   current $\Delta cost+\hat{cost}_{r}([P_{i}]_{cut^{*}})<cost\_limit$ then
  Recluster($P_{i}]_{cut^{*}}$)
Additionally, we dynamically adjust the sliding window size based on its prediction accuracy to adapt to the evolution of the query workload and the data distribution. After completing a window $W$ of queries, we compare the actual cost reduction $\Delta cost_{q}=\sum_{Q\in W}\Delta cost_{q}(Q)$ against the predicted cost reduction $\Delta\hat{cost}_{q}=\sum_{Q\in W}\Delta\hat{cost}_{q}(Q)$. If $\Delta cost_{q}>\Delta\hat{cost}_{q}$, it indicates that the workload is relatively stable, and we therefore double the window size to capture a longer query history. Otherwise, we halve the window size to focus more narrowly on the most recent workload shifts and distribution changes.
@@ 4.3. Cost Model Extensions
where $|\Delta\hat{pruned}(\mathcal{P}\mid Q)|$ denotes the estimated increase in skipped micro-partitions for $Q$ after reclustering $\mathcal{P}$. By tuning $\beta$, users can express their specific exchange rates between cost budget and query performance.
@@ 4.4. Hybrid Layout
To understand how each micro-partition might benefit from reclustering, we decompose its aggregate estimated savings into per-column components. We assume that each predicated column contributes equally in a query. Specifically, for each selected micro-partition $P$, we distribute its aggregated savings across all columns that appear in a query’s predicate within the sliding window:
$s_{P}=\bigl(s_{P}^{(1)},s_{P}^{(2)},\dots,s_{P}^{(d)}\bigr),$
where $d$ is the number of attributes, $s_{P}^{(i)}$ is the share of $P$’s projected benefit attributed to column $i$. Each savings vector is then plotted in a high-dimensional space to create the micro-partition’s “savings signature”, as illustrated in Figure 8.
@@
  • Single-column dominance. If a micro-partition’s projected savings are contributed almost entirely by a single column, reclustering using the column as the clustering key is sufficient. Micro-partitions within the same cluster are grouped and sorted together by their shared dominant column.
@@
The candidate micro-partitions are therefore divided into mutually exclusive groups, each assigned the physical layout that promises the greatest benefit. This strategy produces a hybrid table organization: different regions of the data are clustered with different keys, yet every reclustering decision is still guided by the same workload-aware cost model described earlier. Micro-partitions thus shift between groups as the workload evolves, allowing the table layout to adapt over time.
@@ 5. RECLUSTERING AS A SERVICE
Figure 9. Service Framework Workflow. A metadata service continuously collects partition metadata, workload execution statistics, and reclustering snapshots. After a query completes, a reclustering operation may be triggered, which uses the metadata to determine its scope and is then executed on a dynamically provisioned compute node.
Service Framework Workflow
Having presented the WAIR algorithm, we now introduce the Automatic Reclustering Service (ARS), a scalable reclustering framework that autonomously manages background operations to preserve table clustering quality over time. ARS leverages execution statistics with minimal compute overhead to decide the scope and timing of reclustering. ARS moves all reclustering work off the query-execution critical path, ensuring zero disruption to production workloads in practice. In this section, we describe how to embed the WAIR algorithm into ARS with minimal engineering effort and margin: full compatibility on top of cloud data warehouses. We present the practical feasibility of ARS and the implementation details to be evaluated, and margin: provide an overview of pruning opportunities across representative query patterns.
@@ 5.1. Service Framework
As shown in Figure 9, integrating the reclustering framework into an existing cloud data warehouse additionally requires: 1) an (often already in place) metadata service, and 2) a dedicated reclustering job executor.
@@
Metadata is continuously tracked and preserved during data ingestion and query execution asynchronously for further utilization. margin: In modern distributed database systems, the metadata service is typically a standalone component. It should integrate seamlessly with a cloud data warehouse’s existing metadata infrastructure as a drop-in module.
Reclustering Executor. After specifying the partition set to be reclustered and corresponding layouts, the reclustering executor retrieves these partitions from storage layer, loads their data into local memory, and reorganizes them into new partitions. These clustered partitions are then persisted back to the object storage.
@@
Reclustering operations are designed to be non-blocking in distributed systems. When a reclustering operation completes, ARS publishes a new table snapshot that references the current set of partitions. Because reclustering reorganizes data without changing any row values, queries that began before completion continue to read the prior snapshot along with any newly ingested partitions. New queries started after completion read the reclustered partitions referenced by the latest, updated snapshot. Once no queries reference an older snapshot, ARS safely removes obsolete partitions during garbage collection. The table snapshot chain incurs minimal storage overhead and allows new queries to benefit immediately once reclustering completes. Reclustering tasks run on dynamically provisioned compute nodes, entirely outside the query execution critical path. margin: In cloud data warehouses with a cache layer, reclustering naturally benefits from scan sharing as it requires scanning complete micro-partition data. Moreover, systems can opt for the k-th quantiles method (Cormen et al., 2022) over standard, well-optimized sorting to recluster partitions in $O(n\log k)$ time.